40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu May 16 04:46:37 2013, Jenne, Update, LSC, Kiwamu's sensing matrix measurement script revived 
    Reply  Thu May 16 23:48:39 2013, Jenne, Update, LSC, Kiwamu's sensing matrix measurement script revived 
       Reply  Mon May 20 17:49:36 2013, Jenne, Update, LSC, PRMI sensing matrix - not high quality data PRMIsensMatQuickAnalysis.m.gz
          Reply  Mon May 20 18:47:47 2013, Koji, Update, LSC, PRMI sensing matrix - not high quality data 
             Reply  Tue May 21 14:48:08 2013, Jenne, Update, LSC, PRMI sensing matrix - not high quality data 
       Reply  Mon May 20 18:50:22 2013, Jenne, Update, LSC, Kiwamu's sensing matrix measurement script revived 
          Reply  Tue May 21 18:22:18 2013, Jenne, Update, LSC, Sensing matrix scripts modified to include actuator calibration 
             Reply  Wed May 22 00:08:19 2013, Koji, Update, LSC, Sensing matrix scripts modified to include actuator calibration 
                Reply  Wed May 22 18:24:19 2013, Jenne, Update, SUS, Violin mode survey 
                Reply  Wed May 22 20:50:26 2013, Jenne, Update, LSC, Sensing matrix scripts don't calculate correctly 
                   Reply  Thu May 23 14:45:08 2013, Jenne, Update, LSC, Sensing matrix scripts calculations make more sense now SensMat_JenneMeas_23May2013.pngSensMat_KojiMeas_23May2013.png
                      Reply  Thu May 23 21:45:51 2013, Jenne, Update, LSC, Sensing matrix scripts now check for lockloss 
                Reply  Thu May 23 19:09:15 2013, Jenne, Update, LSC, Sensing matrix scripts modified to include actuator calibration SensMat_21May2013.png
          Reply  Wed May 22 18:07:36 2013, Jenne, Update, LSC, Kiwamu's sensing matrix measurement script revived 
             Reply  Thu May 23 00:49:13 2013, Jenne, Update, LSC, LSC filters loaded 
Message ID: 8603     Entry time: Tue May 21 14:48:08 2013     In reply to: 8601
Author: Jenne 
Type: Update 
Category: LSC 
Subject: PRMI sensing matrix - not high quality data 

The PRMI sensing matrix, as measured last Thursday, in a more readable format:

EDIT: DON'T Look at this yet!  I forgot to calibrate it!  Please hold.....

            PRCL          MICH        
AS55_I      1.228E-01     5.476E-03    
AS55_Q      1.547E-01     1.345E-01    
REFL11_I    9.946E+03     8.106E+01    
REFL11_Q    2.346E+03     2.707E+01    
REFL33_I    9.639E+01     8.717E-01    
REFL33_Q    9.707E-01     6.626E-02    
REFL55_I    1.040E+02     8.464E-01    
REFL55_Q    2.018E+00     5.803E-01

 

Okay, Calibrated, but forgot to include loop compensation (since notches didn't exist yet):

Sensing Matrix, units = cts/meter
 
            MICH          PRCL        
AS55_I      5.024E+08     9.418E+07    
AS55_Q      6.328E+08     2.313E+09    
REFL11_I    4.068E+13     1.394E+12    
REFL11_Q    9.594E+12     4.656E+11    
REFL33_I    3.942E+11     1.499E+10    
REFL33_Q    3.970E+09     1.140E+09    
REFL55_I    4.253E+11     1.456E+10    
REFL55_Q    8.252E+09     9.981E+09

 

ELOG V3.1.3-